A particle of mass 5.0 × 10–3 kg, moving with simple harmonic motion of amplitude 0.15 m, takes 47 s to make 50 oscillations. What is the maximum kinetic energy of the particle?
Answer #1

Max kinetic energy of particle is given by:

KEmax = 0.5*m*Vmax^2

Vmax = A*w

A = Amplitude = 0.15 m

w = 2*pi/T

T = time period = 47 s/50 oscillation = 0.94 sec

So,

Vmax = 0.15*2*pi/0.94 = 1.00 m/sec

m = mass of particle = 5*10^-3 kg

So,

KEmax = 0.5*5*10^-3*1.00^2

KEmax = 2.5*10^-3 J

Please Upvote.

Comment below if you have any query.